OpenStudy (anonymous):

Help please!! What are the real and complex solutions of the polynomial equation? x^4-41x^2=-400

OpenStudy (anonymous):

t=x^2 t^2-41t+400=0 t=16 or t=25 if t=16 then x^2=16 x=4 or x=-4 if t=25 then x^2=25 x=5 or x=-5 Result: x=+-4 or +-5 - all solutions are real.

OpenStudy (michele_laino):

the complete factor of your polynomial is: \[(x ^{2}-25)*(x ^{2}-16)\] so you have to solve two quadratic equations

OpenStudy (michele_laino):

the answer of @AndreyZn is a good answer!